OpenStudy (anonymous):

Does an inverse exist for the following? |-2 1| |-6 1|

OpenStudy (anonymous):

Use the simple trick for 2X2 matrices \[\frac{ 1 }{ ad-bc }\ \left[\begin{matrix}-d & b \\ a & -c\end{matrix}\right]\]

OpenStudy (anonymous):

an inverse exists if the determinant is not = 0

OpenStudy (anonymous):

determinant being ad - bc

OpenStudy (anonymous):

so the latex command for making a matrix is: \left[\begin{matrix}-d & b \\ a & -c\end{matrix}\right] good to know =D ok so I got \[\frac14\left[\begin{matrix}1 & -6 \\ 1 & -2\end{matrix}\right]\] since that gives me a value.....It has an inverse!!!!! Thanks guys!!!